OpenStudy (calculusxy):

If A = (3, 5 + t) and B = (2t, 6), find the vectors ->(AB) and ->(BA)

OpenStudy (calculusxy):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

vector AB means we start at A and move to B A is the tail B is the tip 

OpenStudy (calculusxy):

->(AB) would be [2t - 3, 1 + t]

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

you subtract `tip - tail`

jimthompson5910 (jim_thompson5910):

wait

jimthompson5910 (jim_thompson5910):

it's not 1+t, it should be 1-t

jimthompson5910 (jim_thompson5910):

6-(5+t) = 6-5-t = 1-t

OpenStudy (calculusxy):

ah yes

OpenStudy (calculusxy):

vector ba it would be [3 - 2t, -1 + t]

jimthompson5910 (jim_thompson5910):

yep BA is correct

OpenStudy (calculusxy):

Thanks again!

jimthompson5910 (jim_thompson5910):

sure thing